These were accepted apart from the following amendment under Actions Completed:
Toby and Neil are happy that non-ASCII names shouldn't be a problem.
has been replaced with Toby and Neil don't anticipate that non-ASCII names will cause problems but will try to fix if required.
A number of sites have been suggested but we need more. Craig noted that he had identified some.
Neil's proposals were accepted. Services Unit to tarball list contents for those to be deleted. For lists that are to remain, list owners should ensure membership is correct.
Alastair would like some examples to help diagnose the problem. US will provide.
US will ask Units about kit and if to be kept, will record reason.
Stephen reckons there will be a 'flush all' option with 1.6.6
Will set retention times on queue by queue basis. Alison to send proposal to be sent to George/Iain. Now being tracked at CEG.
Have emailed Angus Rae asking more generally how the CO community communicate.
Alastair further clarified the situation
Stephen/Craig/Alison to confirm documentation up to date including a page listing instructions by OS.
IS responded to say that they intend implementing a random password generator. Using EASE is not a simple option.
It was agreed that this isn't a problem. External examiners very rarely see the exam scripts until BoE meetings.
Blog articles completed:
Report from Computing Executive Group
Reports from units
In summary, it was agreed that US will take responsibility for harvesting the results and Services will be responsible for deciding on appropriate action including advising users with self-managed machines.
Please ensure cardboard boxes are moved away from beside the Physics DR rack.
Alison noted that this has uncovered a few issues, particularly with visitor accounts. The problem mainly lies with DICE accounts (and information in Theon) not being in sync with details held centrally and will give some thought to improving this.
Discussions on SL6.4 upgrades came to an agreement that the DICE_STICK_WITH_SL63 macro would be ignored after 28th August
Attention was drawn to the possible confusion with firmware upgrades reported on ordershost. See MP report for details.
Topics for discussion
We went through the meeting notes. C(S)O's were reminded to upgrade their own home and self-managed machines, and to report how things go. We seem to be on track to meet the deadline of Friday 30th August for server reboots and the subsequent dropping of DES on 2nd September.
Windows machines will be upgraded to 1.7.26 manually.
Craig/Alison/Stephen will liaise on ensuring documentation is updated. Craig has written a blog article.
Wednesday 28th August at 10:00 in IF-4.31
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh